Murphy Inc., which produces a single product, has provided the following data for its most recent month of operation:

Number of units produced 7,000
Variable costs per unit:
Direct materials $ 86
Direct labor $ 71
Variable manufacturing overhead $ 2
Variable selling and administrative expenses $ 7
Fixed costs:
Fixed manufacturing overhead $ 378,000
Fixed selling and administrative expenses $ 98,000

The company had no beginning or ending inventories.

Required:

a. Compute the unit product cost under absorption costing.

b. Compute the unit product cost under variable costing.
Answer #1

Fixed manufacturing overhead per unit=(378,000/7000)=$54 per unit

a.Unit product cost=Direct materials+Direct labor+Variable manufacturing overhead+Fixed manufacturing overhead

=(86+71+2+54)=$213

b.Unit product cost=Direct materials+Direct labor+Variable manufacturing overhead

=(86+71+2)=$159